# DHGS: Decoupled Hybrid Gaussian Splatting for Driving Scene

### v1

@@ Introduction
Scene Reconstruction on Waymo Dataset. For reconstruction views, we calculate the mean PSNR, SSIM, and LPIPS of the rendered images for 32 clips as shown in Table 1. Our method outperforms all the mainstream methods: including 3DGS, 2DGS, Gaussian Pro and Scaffold GS. The left part of Figure 4 shows the rendered images of the reconstruction views from our method and the comparison methods. It can be noticed that our method achieves better reconstruction quality with finer details than competitory methods, from which the road area rendered by our method contains richer texture information.
@@
Novel View Synthesis on Waymo dataset. We evaluate the aforementioned methods on the novel views. Table 1 illustrates the validation results, our method surpasses most rival methods, such as 3DGS, 2DGS and Gaussian Pro, and achieves comparable results to Scaffold GS. Here Scaffold GS gets the best performance on PSNR and SSIM. As depicted in the right part of Figure 4, blurriness arises in some non-road areas (framed in red boxes) for 3DGS, 2DGS, and Gaussian Pro. Meanwhile, our method restores more road details (lane lines within the red boxes) due to the imposition of geometric constraints on the road, whereas other methods such as 2DGS, Gaussian Pro, and Scaffold GS fail to retain these details.
@@
Free-view Novel View Synthesis on Waymo dataset. Although the novel views are not utilized for training, the changes in their camera poses are relatively small compared to the trained views, along with no variation in the relative position of the cameras to the vehicle, which can be simply concluded as the interpolation of trained views. To further demonstrate the free-view novel view synthesis performance of our method, evaluating the trained model on greater pose variations of cameras, we vary the poses of the cameras relative to the vehicle, simulating various observation perspectives for different vehicle types with diverse camera positions. In this paper, we perform four transformations on the camera poses, the specific configurations are detailed in the supplemental material. Since there is no ground truth for the free-view novel views, we adopt FID metric that is able to effectively measure the feature gap between the distributions of the origin input views and the free-view novel views. The comparisons of FID for free-view novel views are also presented in Table 1, where our method achieves the highest performance compared to other methods in all four configurations with considerable preponderance. The qualitative render results are shown in Figure 5, where the free-view rendered images produced by other methods suffer from extensive artifacts and lane line breakage. Our method is still able to render complete road details such as lane lines and environmental details even after significant changes in the camera pose. The above results further confirm that the regularization constraints we proposed can ensure the model learns a more optimal geometric road structure, thereby showcasing the robust capabilities of our approach in untrammeled novel view synthesis. More visual comparisons can be found in supplemental materials.
@@
Figure 6: Qualitative comparison for free-view novel views of ablation results. The hightlighted patches demonstrate that these regularization terms are all beneficial to the reconstruction of road geometry.
@@
To validate the effectiveness of the proposed regularization components, we conduct an ablation study. We train our proposed method by separately removing the SDF loss, transmittance loss and consistency loss, denoted as ”w/o SDF”, ”w/o trans”, and ”w/o cons” respectively. The quantitative indexes of scene reconstruction, novel view synthesis and free-view novel view synthesis are presented in Table 2. For scene reconstruction and novel view synthesis, removing the SDF loss leads to a decline in various metrics, while removing the transmittance loss and consistency loss does not result in significant drop. However, for free-view novel view synthesis, removing any of the losses leads to an increase in the FID score, indicating a larger discrepancy between the novel views and the original images. To more intuitively illustrate the specific contributions of these components, we give the visualization results of free-view novel views synthesis in Figure 6. When we remove the SDF regularization term, the traffic sign characters on the road undergo noticeable distortions. When the consistency regularization term is removed, a clear dark seam emerges between the road and the environment. If the transmittance regularization term is removed, large areas of artifacts appear on the road. These regularization terms each play their own role. Specifically, the SDF loss ensures the correctness of road geometry, the consistency loss guarantees a seamless connection between the environment model and the road model. The transmittance Loss ensures that the rendering areas of the road model and the environment model accurately align with the 2D semantic segmentation map, thereby preventing the occurrence of artifacts when rendering novel views.
@@
Table 2: Quantitative results of ablation experiments on Waymo dataset, with each metric averaged across all clips.
@@ Conclusions
In this paper, we propose a novel reconstruction method called DHGS, which aims to enhance the rendering quality of novel view synthesis for driving scenes. The method uses two models to model the road and environment respectively, and introduces a pixel-level blending renderer that overlays the rendered layers of these two models through the proposed depth-ordered rendering strategy. Additionally, consistency loss and transmittance loss are proposed to constrain the fundamental models, enabling the generation of continuous and consistent rendered images. Furthermore, we propose a road surface regularization term based on SDF to ensure better geometric consistency of the road surface, guaranteeing stable image quality as perspective changes. We conduct comprehensive experiments on the Waymo dataset, revealing the state-of-the-art performance in both reconstruction and novel view synthesis of our method.
@@ Acknowledgments
This work would thank Bo Zheng and Fan Jin for providing necessary help in the preparation of training data.
Supplemental Material for DHGS: Decoupled Hybrid Gaussian Splatting for Driving Scene
@@ A. Implementation Details
To deal with the floaters phenomenon in new perspectives, we devise a simple yet effective strategy. We construct a hemispherical virtual sky point cloud layer surrounding the LiDAR point clouds. The radius and center of this sky point cloud were calculated based on the scale of the and the average distribution of the LiDAR points, aiming to accurately mimic the spatial properties of the sky. Additionally, we apply appropriate coloring to the sky point cloud to ensure its visual fidelity, which is shown in Figure 7. For fairness, we integrate this sky sphere point cloud mechanism into all comparative methods.
Figure 7: Visualization of incorporating a sky sphere point cloud to the initial point cloud.
@@
Figure 8: The simplified sketch of distance constraint and normal constraint guided by the zero-level set of SDF, the function $g$ refers to the ideal pretrained SDF.
@@
The SDF regularizer $\mathcal{L}_{\text{sdf}}$, inspired by NeuSG originally optimizing the positions and orientations of 3D Gaussians, is applied in this paper similarly to 2D Gaussians. Specifically, we leverage the gradient direction of the pre-trained SDF to guide the orientation updates of the 2D Gaussians, while simultaneously constraining their disk centers to align with the SDF. This approach ensures a seamless integration which is intuitively illustrated in Figure 8.
@@
Table 3: The table illustrates four different parameter settings for free-view novel view synthesis, where Translation refers to translation component, Rotation refers to rotation component, the corresponding three numbers are the lengths of translation in meter and the magnitudes of three Euler angles for rotation in degree.
@@ B. Experimental results
Free-view Noval View Synthesis Settings.
@@
We conduct four transformations to the pose of the camera relative to the vehicle, including rotations and translations. The detailed variation amplitudes for the front, front-left, and front-right camera under each transformation are presented in Table 3, where Translation denotes the translation distances of the camera along the $x$, $y$, and $z$ axes in meter, and Rotation represents the rotation angles along the $yaw$, $pitch$, and $roll$ axes in degree. Figure 9 depicts the differences in the four transformed camera poses compared to the original Waymo setup.
Figure 9: The camera layouts after applying the four settings of camera pose. The top row represents the camera layouts after applying various camera pose settings from a bird’s eye view, while the bottom row shows the same transformations from a side view.
@@
Figure 10: Qualitative comparisons on the Waymo dataset, the highlighted regions demonstrate that our method achieves the best rendering quality for both scene reconstruction in the left column and novel view synthesis in the right column.
Qualitative result on Waymo dataset.
@@
In addition to the quantitative metrics elaborated in the main document, we attach additional qualitative results to provide a more intuitive performance evaluation of each method. As shown in Figures 10, 11, and 12, the qualitative results indicate that our method achieves the best visualization result in the modeling of both environmental backgrounds and road structures, demonstrating superior performance in scene reconstruction, novel view synthesis and free-view novel view synthesis.
@@
Figure 11: Qualitative results of free-view novel view synthesis under different scenarios, with the left column presenting the outputs from Set3 and the right column showing those from Set4. The highlighted red boxes demonstrate that our method excels in capturing fine details both in terms of geometry and appearance.
@@
Figure 12: Qualitative results of free-view novel view synthesis under different scenarios, with the left column presenting the outputs from Set3 and the right column showing those from Set4. The highlighted red boxes demonstrate that our method excels in capturing fine details both in terms of geometry and appearance.

### v2

@@ Introduction
Scene Reconstruction on the Waymo Dataset. For reconstruction views, we calculate the mean PSNR, SSIM, and LPIPS of the rendered images across 32 clips, as shown in Table 1. Our method outperforms all the mainstream methods: including 3DGS, 2DGS, Gaussian Pro, and Scaffold GS. The left part of Figure 4 displays the rendered images of the reconstruction views from our method and its competitors. It can be noticed that our method achieves better reconstruction quality with finer details than competitive methods, from which the road area rendered by our method contains richer texture information.
@@
Novel View Synthesis on the Waymo dataset. We assess the aforementioned methods on the novel views. Table 1 illustrates the validation results, demonstrating that our method surpasses most rival methods, such as 3DGS, 2DGS, and Gaussian Pro, and achieves comparable results to Scaffold GS. Notably, Scaffold GS achieves the best performance in terms of PSNR and SSIM. As depicted in the right part of Figure 4, blurriness appears in some non-road areas (highlighted by red boxes) for 3DGS, 2DGS, and Gaussian Pro. In contrast, our method restores more road details (lane lines within the red boxes) thanks to the imposition of geometric constraints on the road, whereas other methods such as 2DGS, Gaussian Pro, and Scaffold GS fail to preserve these subtleties. Despite novel views not being part of the training set, their camera poses variations are minimal compared to trained views, which can be simply concluded as the interpolation of trained views. Such minor variations in novel views do not sufficiently reflect the geometric representation capability of the model.
@@
Free-view Novel View Synthesis on the Waymo dataset. To further underscore the geometric consistency of our approach, we evaluate the trained model on free-view novel views with more pronounced camera pose variations. We adjust the poses of the cameras relative to the vehicle, mimicking various observation perspectives for different vehicle types with diverse camera positions. In this paper, we outline four types of transformations to the camera poses, with detailed configurations provided in the supplementary material. Given the absence of ground truth for these novel views, we adopt the FID metric, adept at gauging the feature discrepancy between the original input views and the free-view novel views. The comparisons of FID for free-view novel views are also presented in Table 1, where our method achieves the highest performance compared to other methods in all four configurations with a significant margin of superiority. The qualitative rendering results, depicted in Figure 5, reveal that free-view rendered images produced by other methods suffer from extensive artifacts and lane line breakage. In contrast, our method consistently captures environmental and road details, including lane lines, even amidst substantial camera pose shifts. The results further validate that the regularization constraints and the depth-ordered hybrid rendering strategy we proposed facilitate the acquisition of an optimal geometric structure, thereby showcasing the robustness of our approach in free-view novel view synthesis. More visual comparisons can be found in supplemental materials.
@@
Figure 6: Qualitative comparison for free-view novel views of ablation studies. The emphasized patches reveal that these regularization terms contribute to the accurate reconstruction of road geometry.
@@
To assess the efficacy of our proposed regularization components, we perform an ablation study, systematically omitting the SDF, transmittance, and consistency losses, denoted as ”w/o SDF”, ”w/o tran”, and ”w/o cons”. The quantitative results for scene reconstruction, novel view synthesis, and free-view novel view synthesis are consolidated in Table 2. For scene reconstruction and novel view synthesis, the absence of the SDF loss leads to a decline in performance metrics, whereas the exclusion of transmittance loss or consistency loss has a minimal impact. Conversely, for free-view novel view synthesis, the removal of any loss escalates the FID score, signifying a greater divergence between the novel views and the originals. To more intuitively illustrate the specific contributions of these components, we present the visualization results of free-view novel views synthesis in Figure 6. Specifically, the SDF loss ensures the fidelity of road geometry, the consistency loss guarantees a seamless connection between the environment model and the road model. The transmittance loss ensures precise alignment of the rendering areas of both models with the 2D segmentation, thus averting rendering artifacts. Therefore, each regularization term is integral to our DHGS framework.
@@
Table 2: Quantitative results of ablation experiments on the Waymo dataset, with each metric averaged across all clips.
@@ Conclusions
In this paper, we propose a novel reconstruction method called DHGS, which aims to elevate the rendering quality of novel view synthesis for static driving scenes. DHGS leverages two decoupled models to represent the road and environment separately, which are then merged via a pixel-level hybrid renderer, facilitated by our innovative depth-ordered hybrid rendering strategy. Additionally, consistency loss and transmittance loss are proposed to constrain the basal models, allowing for the generation of continuous and consistent rendered images. Moreover, we develop a road surface regularization term grounded in SDF to ensure better geometric consistency of the road surface, guaranteeing stable image quality as perspective changes drastically. Through extensive testing on the Waymo dataset, our method reveals the state-of-the-art performance in both reconstruction and novel view synthesis of our method.
@@ Acknowledgments
This work would thank Bo Zheng and Fan Jin for providing the necessary help in the preparation of training data.
Appendix A Supplemental Material for DHGS: Decoupled Hybrid Gaussian Splatting for Driving Scene
@@ A. Implementation Details
To deal with the floaters phenomenon in new perspectives, we devise a straightforward yet effective strategy. We construct a hemispherical virtual sky point cloud layer that envelops the LiDAR point clouds. The radius and center of this sky point cloud are determined by the scale and the average distribution of the LiDAR points, ensuring an accurate representation of the sky’s spatial characteristics. Additionally, we apply appropriate coloring to the sky point cloud to maintain its visual integrity, as depicted in Figure A1. Table A1 presents a comparative analysis between incorporating the sky sphere point cloud into 3DGS and not incorporating it. It is evident that the introduction of sky sphere pcd is able to improve the rendering quality, no matter in reconstruction or novel view synthesis. For fairness, we integrate this sky sphere point cloud mechanism into all comparative methods, such that the potential biases stemming from differences in sky handling can be eliminated.
Figure A1: Visualization of incorporating a sky sphere point cloud into the initial point cloud.
{tblr}
[…]
PCD wo sky sphere 25.63 0.8062 0.3265 67.52
PCD wi sky sphere 25.92 0.8089 0.3218 64.77
Table A1: The quantitative results of 3DGS with and without incorporating the sky sphere point cloud into the initial point cloud. The FID here is averaged on FIDS of all four sets.
@@
Figure A2: The simplified sketch of the SDF regularizer, where $\mathcal{L}_{\text{dist}}$ and $\mathcal{L}_{\text{norm}}$ refer to the distance constraint and normal constraint guided by the zero-level set of SDF, the function $g$ refers to the ideal pretrained SDF.
@@
The SDF regularizer $\mathcal{L}_{\text{sdf}}$, inspired by NeuSG originally optimizing the positions and orientations of 3D Gaussians, is applied in this paper similarly to 2D Gaussians. Specifically, we leverage the gradient direction of the pre-trained SDF to guide the orientation updates of the 2D Gaussians, while simultaneously constraining their disk centers to align with the SDF. This approach guarantees a seamless integration which is intuitively illustrated in Figure A2.
@@
Table A2: The table illustrates four different parameter settings for free-view novel view synthesis, where Translation refers to the translation component, Rotation refers to the rotation component, the corresponding three numbers are the lengths of translation in meter, and the magnitudes of three Euler angles for rotation in degree.
Figure A3: The camera layouts after applying the four settings of camera pose. The top row represents the camera layouts after applying various camera pose settings from a bird’s eye view, while the bottom row shows the same transformations from a side view.
@@ B. Experimental results
Free-view Novel View Synthesis Settings.
@@
We apply four transformations to the pose of the camera relative to the vehicle, including rotations and translations. The motive is that transferring the visual data from the collecting vehicle to a new vehicle with a diverse sensor setup is frequently demanded in AD. We set four settings based on the camera positions of the Waymo dataset to simulate the camera displacements caused by changes in vehicle type. The detailed variation amplitudes for the front, front-left, and front-right camera under each transformation are outlined in Table A2, where Translation denotes the translation distances of the camera along the $x$, $y$, and $z$ axes in meter, and Rotation represents the rotation angles arorund the $yaw$, $pitch$, and $roll$ axes in degree. Figure A3 depicts the differences in the four transformed camera poses compared to the original Waymo configuration.
{tblr}
[…]
Train Test Set1 Set2 Set3 Set4
$\text{All}_{\text{3DGS}}$ 26.91 25.92 53.12 61.93 82.52 61.53
$\text{All}_{\text{Scaffold GS}}$ 28.04 26.87 51.93 58.25 78.64 57.85
$\text{E}_{\text{3DGS}}$ + $\text{R}_{\text{3DGS}}$ 26.86 25.93 52.90 60.38 80.49 58.14
$\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{3DGS}}$ 27.91 26.47 48.95 58.69 77.92 57.41
$\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{2DGS}}$ 28.09 26.77 48.05 56.94 75.72 55.52
[…]
@@
Figure A4: Qualitative comparisons of different compound modes on the Waymo dataset, with top ($\text{E}_{\text{3DGS}}$ + $\text{R}_{\text{3DGS}}$), middle ($\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{3DGS}}$), and bottom ($\text{E}_{\text{Scaffold GS}}$ + $\text{R}_{\text{2DGS}}$). The highlighted regions demonstrate that our method achieves the best rendering quality.
Figure A5: Qualitative comparisons on the Waymo dataset, the highlighted regions demonstrate that our method achieves the best rendering quality for both scene reconstruction in the left column and novel view synthesis in the right column.
The qualitative result on the Waymo dataset.
@@
In addition to the quantitative metrics and qualitative rendered images elaborated in the main document, we attach additional qualitative results to provide a more intuitive performance evaluation of each method. As shown in Figures A5, A6, and A7, the qualitative results indicate that our method achieves the best visual quality in the modeling of both environmental backgrounds and road structures, demonstrating superior capability in scene reconstruction, novel view synthesis, and free-view novel view synthesis.
Comparative analysis of the compound mode of Gaussians.
We compare several distinct assembly ways of road and non-road models, as detailed in Table A3. Figure A4 illustrates qualitative comparisons of three compound modes. It can be concluded that leveraging Scaffold GS and 2DGS to represent the environment and road model severally harvests optimal performance. Benefiting from the robustness of Scaffold GS in changing views and the inherent superiority of flat 2DGS representing the road surface, DHGS achieves superior rendering results both in the road and environment areas, exhibiting fewer floaters on the road and capturing finer environmental details..
@@
Figure A6: Qualitative results of free-view novel view synthesis under different scenarios, with the left column presenting the outputs from Set3 and the right column showing those from Set4. The highlighted red boxes demonstrate that our method excels in capturing fine details both in terms of geometry and appearance.
@@
Figure A7: Qualitative results of free-view novel view synthesis under different scenarios, with the left column presenting the outputs from Set3 and the right column showing those from Set4. The highlighted red boxes demonstrate that our method excels in capturing fine details both in terms of geometry and appearance.